OpenStudy (anonymous):

solve the equation 7y-3y-14=14

OpenStudy (anonymous):

combine 7y - 3y add 14 to both sides divide y to both sides..

OpenStudy (anonymous):

7y-3y=4y 14+14=28 28/4=7

OpenStudy (anonymous):

y=7